OpenStudy (anonymous):

The line of symmetry for the quadratic equation y = ax^2 - 2x - 3 is x = 1. What is the value of "a"?

OpenStudy (anonymous):

-b/2a =1

OpenStudy (anonymous):

b=-2 -b=2 2/2a=1 2=2a a=1